Solve 2(x-1)^2 = -16 over the set of complex numbers.
Answer by Nate(3500)

2(x - 1)^2 = -16
(x - 1)^2 = -8
x - 1 = +-sqrt(-8)
x = 1 +- 2i*sqrt(2)